Creating a mailing

Johannes Braams

september 7, 1994

1 Introduction

This package is intended to be used when you want to send a large number of letters, all with (almost) the same text.

This package is based on the former style option merge by Graeme McKinstry, but is is a reimplementation with a different user interface.

2 The user interface

\addressfile The commands \addressfile takes a filename as its argument. When the file

can’t be found by L^{A}TEX, the user will be asked to supply a different name. The address file should have the following format:

Name of addressee
Street\\town
Opening text of the letter
(optional definitions)
<blank line>
Name of addressee
Street\\town
Opening text of the letter
(optional definitions)
<blank line>
...

The various addresses are separated by a blank line in the file (multiple blank lines in between addresses are allowed). It is also possible to have multiple lines with definitions; they will all be executed.

\mailingtext The text of the mailing is entered as the argument of \mailingtext. A difference with the original merge.sty is that this package allows control sequences in the argument of \mailingtext. These control sequences should then be defined in the file with the address information.

\makemailing When \makemailing is called the letters are produced, combining the information found in the address file with the text of the mailing.